# Learning a Unified Degradation-aware Representation Model for Multi-modal Image Fusion

### v2

@@ 4.4.1 Qualitative Comparison
Unlike the existing all-in-One method (DAFusion [48]), LURE benefits from explicit labels as supervisory signals, resulting in richer textural details and less noise in fusion outcomes. Compared to Text-IF [56], our method unconstrained by quadruple data format, leverages more high-quality image restoration datasets with real-world scenarios, and yielding more natural images without pre-enhanced sources. Against two-step models like CSCFuse [65] or MMDRFuse [11], LURE reduces domain shift-related degradation and information loss.
More crucially, consistent with DAFusion, our method inherently addresses cross-modal combined degradations, achieving superior performance without supplementary combined degradation training datasets. Text-IF, method-limited, struggles with combined degradations, thus needs external restoration, yielding lower quality. Specifically, for OE+LC and LL+LC tasks, our method exhibits superior texture detail, color vividness, and contrast compared to other approaches. This highlights our method’s ability to attain high-quality fused images without additional restoration models or even combined degradation training datasets.
@@ 4.5 Multi-modal Semantic Segmentation
To assess the effectiveness of LURE in high-level tasks, the experiment of multi-modal semantic segmentation is conducted. Segformer-b2 [53] is fine-tuned on MSRS fusion results and evaluated on the test dataset  [^4] For more details please refer to supplementary material..
@@
Table 5: Quantitative comparison of multi-modal semantic segmentation on MSRS. Bold/underlined values: best/second best results.
IR VI DCI CSC MMD GTM FuB TIF DAF LURE
mIOU 77.679 78.94 80.05 77.2 78.87 79.4 78.84 78.21 79.1 81.22
@@ 4.6 Ablation Study
Table 6: Ablation Studies on Low Contrast Enhancement (LC), Overexposure (OE), and Low Light Enhancement and Contrast Enhancement for infrared and visible images tasks respectively (LL+LC). Bold/underlined values: best/second best results.
@@
Ours 0.170 0.141 4.621 0.212 1.374 0.251 7.422 55.403 0.150
@@ 5 Conclusion
This paper addresses limitations of prior ADFMs constrained by long-tailed data, dimensionality curse, and synthetic data reliance. By decoupling modality and quality dimensions at data level and reassociating it at ULFS, our proposed approach mitigates these issues, provides effective supervision, and yields superior fusion performance. Furthermore, an inner residual model and Text-Guided Attention (TGA) enhance spatial perception and detail preservation. Extensive experiments validate our proposed method achieves better performance in vanilla and degradation-aware fusion. Importantly, our method is applicable not only to infrared-visible image fusion but also to other multi-modal image fusion tasks. We believe our approach offers a fresh perspective and inspires future fusion research.

### v3

@@ 4.4.1 Qualitative Comparison
Unlike the existing all-in-One method (DAFusion [wang2025degradation]), LURE benefits from explicit labels as supervisory signals, resulting in richer textural details and less noise in fusion outcomes. Compared to Text-IF [yi2024text], our method unconstrained by quadruple data format, leverages more high-quality image restoration datasets with real-world scenarios, and yielding more natural images without pre-enhanced sources. Against two-step models like SAGE [wu2025every] or MMDRFuse [deng2024mmdrfuse], LURE reduces domain shift-related degradation and information loss.
Crucially, consistent with DAFusion, our method inherently addresses cross-modal combined degradations, achieving superior performance without additional combined degradation training datasets. Text-IF, method-limited, struggles with combined degradations, thus needs external restoration, yielding lower quality. Specifically, for OE+LC and LL+LC tasks, our method exhibits superior detail, color vividness, and contrast compared to other approaches. This shows our method can produce high-quality fused images without extra restoration models.
@@ 4.5 Multi-modal Semantic Segmentation
To assess the effectiveness of LURE in high-level tasks, the experiment of multi-modal semantic segmentation is conducted. Segformer-b2 [xie2021segformer] is fine-tuned on MSRS fusion results and evaluated on the test dataset.  [^4] For more details please refer to supplementary material.
@@
Table 5: Quantitative comparison of multi-modal semantic segmentation on MSRS.
IR VI LRRN CDDF MMDR SAGE FuB Text TITA Ours
mIOU 63.09 66.81 66.72 67.38 69.11 68.64 67.33 66.26 68.37 69.25
@@ 4.6 Ablation Study
Table 6: Ablation experiment results.
@@
Ours 0.170 0.150 4.671 0.212 1.374 0.259 7.422 55.403 0.152
Figure 7: Degradation classification accuracy of three classifiers (Input, GT, ULFS) on IR and VI modalities.
4.7 Validation of Latent Representations
To verify whether ULFS is successfully learned, we train three ResNet-based classifiers using degraded inputs, ground-truth images, and ULFS features from various restoration datasets, respectively. The inputs contain degradation cues, the ground truths retain only dataset-specific scene distributions, while ULFS features are expected to exclude degradation and preserve scene information.
We evaluate all classifiers using ACC. As shown in Fig.7, for both IR and VI, the ULFS-based classifier attains accuracy close to the ground-truth one, indicating that degradation information has been effectively removed and the ULFS space has been effectively approximated.
@@ 5 Conclusion
This paper addresses limitations of prior ADFMs constrained by long-tailed data, dimensionality curse, and synthetic data reliance. By decoupling modality and quality dimensions at data level and reassociating it at ULFS, our proposed approach mitigates these issues, provides effective supervision, and yields superior fusion performance. Furthermore, an inner residual model enhance spatial perception and detail preservation. Extensive experiments validate our proposed method achieves better performance in vanilla and degradation-aware fusion. Importantly, our method is applicable not only to infrared-visible image fusion but also to other multi-modal image fusion tasks. We believe our approach offers a fresh perspective and inspires future fusion research.